# Conformally rigid graphs

### v1

@@ 5.1. The complete graph
We give a simple proof that the complete graph is conformally rigid. This proof is the only proof we have that where we show that the constant weights are the only choice of weights maximizing $\lambda_{2}$ and minimizing $\lambda_{n}$; it is structurally quite different from the other proofs in the paper and makes heavy use of the particular structure of the complete graph.
@@ 5.3. Distance-Regular Graphs
which is indeed a nonnegative constant independent of $i$ and $j$. It remains to show that it is a positive constant. We argue by contradiction. If the constant was 0, then we would have $u_{i}=u_{j}$ for all $(i,j)\in E$. Since the graph is connected, this means that $u_{i}=u_{j}$ for all $1\leq i,j\leq n$. This contradicts the fact that the columns of $U$ are an orthonormal set as soon as there are at least two columns. So, suppose $U\in\mathbb{R}^{n\times 1}$ which happens if the dimension of the eigenspace $\mathcal{E}_{\lambda}$ is 1. Appealing again to Lemma 5.3, we see that all entries of $U$ are ±$\pm\sqrt{1/n}$. Since the the column $U$ is an eigenvector of the graph Laplacian and $\lambda>0$, orthogonality to the constant vector implies that both numbers ±$\pm\sqrt{1/n}$ have to appear (and, in fact, have to appear the same number of times), which forces $n$ to be even. This suggests the decomposition $V=[n]=A\cup B$ where
@@ 5.4. Cayley graphs
Suppose there exists an assignment of weights such that $\lambda_{2}(w)>\lambda_{2}(\mathbbm{1})$. This means that for every function $f:G\rightarrow\mathbb{R}$ with mean value 0 and normalized in $\ell^{2}(\Gamma)$, we have
@@
$\sum_{g\in G}(\varphi(g)-\varphi(g\circ s))^{2}$ $=\sum_{g\in G}\varphi(g)^{2}+\varphi(g\circ s)^{2}-2\varphi(g)\varphi(g\circ s))$
@@
Our goal will be to show that for a small value of $\varepsilon$ we can ensure that $\lambda_{2}$ increases. However, it is clear that some restriction on $n$ will be required. When $n=5$, then $C_{5}\left(\left\{1,2\right\}\right)=K_{5}$ is conformally rigid. Moreover, $C_{6}\left(\left\{1,2\right\}\right)$ is the octahedral graph which is conformally rigid (perhaps because it is distance-regular). In both of these cases, the procedure cannot work and it is easy to see from the spectrum why: $K_{5}$ has eigenvalues $(0,5,5,5,5)$ and a perturbation induced by $\varepsilon$ affects all eigenvalues and some will increase while others decrease. Something similar is happening when $n=6$, the octahedral graph has spectrum $(0,4,4,4,6,6)$, in particular, the multiplicity of the second eigenvalue is 3, this allows for it to be conformally rigid.
@@
It suffices to prove the result when $\varepsilon=0$ and then appeal to continuity of the expression in $\varepsilon$. Note first that the expression (32) is the same for $j$ and $-j$, so it suffices to understand the behavior when $j\leq n/2$. We introduce the function
@@ 6. Certifying Conformal Rigidity via SDPs
    Convert the coefficients into reasonable nearby numbers.
@@
This method also worked for the $(20,8)-$accordion graph; the solution $X$ for $\lambda_{n}$ is trivial, the solution $Y$ for $\lambda_{2}$ is not. The numerical solution returned by the solver is not very insightful, however, projecting into the eigenspace corresponding to $\lambda_{2}$ is very helpful because it only has multiplicity 2: the new coordinates happened to either be 0 or, in case they were both nonzero, happen to have a ratio of $(1+\sqrt{5})/2$ which helped identify the correct solution.
@@ 6.3. Symmetry Reduction
Figure 11. Haar 565 (left) and its edge orbits (middle and right).
@@
A more involved example is the distance-$2$ graph of the Klein graph which has $n=24$ vertices and $168$ edges that come in $2$ orbits each of size $84$. In this case, $\lambda_{2}(\mathbbm{1})=11.3542$ and $\lambda_{n}(\mathbbm{1})=16.6458$, each with multiplicity $8$. The eigenvalues of $w_{1}L_{1}+w_{2}L_{2}-I+\frac{1}{24}J$ are:
@@ 6.4. Complementary Slackness
Now suppose $X\succeq 0$ and $LX=\lambda_{2}X$. Then $\operatorname{rank}(X)\leq k$ where $k=\textup{mult}(\lambda_{2})$. Since $X\succeq 0$, $X$ has a reduced SVD of the form $X=ADA^{\top}$ where the columns of $A$ are orthonormal and $D\in{\mathbb{R}}^{k\times k}$ is a diagonal matrix with nonnegative diagonal entries. The column space of $X$ is the column space of $A$ which means that $A=UB$ for some matrix $B\in{\mathbb{R}}^{k\times k}$. Therefore,
@@ 6.5. Using $UU^{\top}$
We conclude by noting a particularly simple method that often ends up producing certificates $X$ and $Y$ (and provably does so for distance-regular graphs). For $\lambda_{2}$, the procedure works as follows: one collects an orthonormal basis of the eigenspace $\lambda_{2}$ corresponding to the Laplacian $L$ in a matrix $U$. A rescaling of the matrix $UU^{\top}\rightarrow cUU^{\top}$ to ensure that the trace is $|E|/\lambda_{2}$ often produces certificates that seem to work: in particular, running a check on the graphs built into the internal database of Mathematica shows that this approach leads to a valid certificate (for $\lambda_{2}$) all conformally rigid graphs having less than 20 vertices. The first case where this approach does not work is the crossing number graph 6B on 20 vertices [24] (see Fig. 1), which also appears in our list of sporadic outliers. Then there are further examples where this approach fails: the Haar graph 565 on 20 vertices (see Fig. 11). However, there are other examples such as the non-Cayley vertex-transitive graph (24,23) on $n=24$ vertices, where all the other approaches failed while $cUU^{\top}$ indeed produces a valid certificate for $\lambda_{2}$.

### v2

@@ 5.1. The complete graph
We give a simple proof that the complete graph is conformally rigid. This proof is the only proof where we show that the constant weights are the only choice of weights maximizing $\lambda_{2}$ and minimizing $\lambda_{n}$; it is structurally quite different from the other proofs in the paper and makes heavy use of the particular structure of the complete graph.
@@ 5.3. Distance-Regular Graphs
which is indeed a nonnegative constant independent of $i$ and $j$. It remains to show that $2m/n-2c_{2}$ is a positive constant. We argue by contradiction. If $2m/n-2c_{2}$ was 00, then we would have $u_{i}=u_{j}$ for all $(i,j)\in E$. Since the graph is connected, this means that $u_{i}=u_{j}$ for all $1\leq i,j\leq n$. This contradicts the fact that the columns of $U$ are an orthonormal set as soon as there are at least two columns. So, suppose $U\in\mathbb{R}^{n\times 1}$ which happens if the dimension of the eigenspace $\mathcal{E}_{\lambda}$ is 1. Appealing again to Lemma 5.3, we see that all entries of $U$ are ±$\pm\sqrt{1/n}$. Since the the column $U$ is an eigenvector of the graph Laplacian and $\lambda>0$, orthogonality to the constant vector implies that both numbers ±$\pm\sqrt{1/n}$ have to appear (and, in fact, have to appear the same number of times), which forces $n$ to be even. This suggests the decomposition $V=[n]=A\cup B$ where
@@ 5.4. Cayley graphs
Suppose there exists an assignment of weights such that $\lambda_{2}(w)>\lambda_{2}(\mathbbm{1})$. This means that for every function $f:G\rightarrow\mathbb{R}$ with mean value 0 and normalized in $\ell^{2}(\Gamma)$, we have, for non-constant weights $w_{g,s}$ on the edge connecting $g$ and $g\circ s$,
@@
$\sum_{g\in G}(\varphi(g)-\varphi(g\circ s))^{2}$ $=\sum_{g\in G}\varphi(g)^{2}+\varphi(g\circ s)^{2}-2\varphi(g)\varphi(g\circ s)$
@@
Our goal will be to show that for a small value of $\varepsilon$ we can ensure that $\lambda_{2}$ increases. However, it is clear that some restriction on $n$ will be required. When $n=5$, then $C_{5}\left(\left\{1,2\right\}\right)=K_{5}$ is conformally rigid. Moreover, $C_{6}\left(\left\{1,2\right\}\right)$ is the octahedral graph which is conformally rigid (it is edge-transitive). In both of these cases, the procedure cannot work and it is easy to see from the spectrum why: $K_{5}$ has eigenvalues $(0,5,5,5,5)$ and a perturbation induced by $\varepsilon$ affects all eigenvalues and some will increase while others decrease. Something similar is happening when $n=6$, the octahedral graph has spectrum $(0,4,4,4,6,6)$, in particular, the multiplicity of the second eigenvalue is 3, this allows for it to be conformally rigid.
@@
It suffices to prove the result when $\varepsilon=0$ and then appeal to continuity of the expression in $\varepsilon$. Note first that the expression (32) is the same for $j$ and $-j$ and the expression is $n-$periodic and thus it is the same for $j$ and $n-j$, so it suffices to understand the behavior when $j\leq n/2$. We introduce the function
@@ 6. Certifying Conformal Rigidity via SDPs
    Convert the coefficients into reasonable nearby numbers (either rational numbers or algebraic numbers of low degree, like $\sqrt{2}$).
@@
This method also worked for the $(20,8)-$accordion graph (see Fig. 7); the solution $X$ for $\lambda_{n}$ is trivial, the solution $Y$ for $\lambda_{2}$ is not. The numerical solution returned by the solver is not very insightful, however, projecting into the eigenspace corresponding to $\lambda_{2}$ is very helpful because it only has multiplicity 2: the new coordinates happened to either be 0 or, in case they were both nonzero, happen to have a ratio of $(1+\sqrt{5})/2$ which helped identify the correct solution.
@@ 6.3. Symmetry Reduction
Figure 11. Haar 565 (left) and its decomposition (middle and right) under the automorphism group of the graph.
@@
A more involved example is the distance-$2$ graph of the Klein graph (see Fig. 7) which has $n=24$ vertices and $168$ edges that come in $2$ orbits each of size $84$. In this case, $\lambda_{2}(\mathbbm{1})=11.3542$ and $\lambda_{n}(\mathbbm{1})=16.6458$, each with multiplicity $8$. The eigenvalues of $w_{1}L_{1}+w_{2}L_{2}-I+\frac{1}{24}J$ are:
@@ 6.4. Complementary Slackness
Now suppose $X\succeq 0$ and $LX=\lambda_{2}X$. Then $\operatorname{rank}(X)\leq k$ where $k=\textup{mult}(\lambda_{2})$. Since $X\succeq 0$, $X$ has a reduced singular value decomposition of the form $X=ADA^{\top}$ where the columns of $A$ are orthonormal and $D\in{\mathbb{R}}^{k\times k}$ is a diagonal matrix with nonnegative diagonal entries. The column space of $X$ is the column space of $A$ which means that $A=UB$ for some matrix $B\in{\mathbb{R}}^{k\times k}$. Therefore,
@@ 6.5. Using $UU^{\top}$
We conclude by noting a particularly simple method that often ends up producing certificates $X$ and $Y$ (and provably does so for distance-regular graphs). For $\lambda_{2}$, the procedure works as follows: one collects an orthonormal basis of the eigenspace $\lambda_{2}$ corresponding to the Laplacian $L$ in a matrix $U$. A rescaling of the matrix $UU^{\top}\rightarrow cUU^{\top}$ to ensure that the trace is $|E|/\lambda_{2}$ often produces certificates that seem to work: in particular, running a check on the graphs built into the internal database of Mathematica shows that this approach leads to a valid certificate (for $\lambda_{2}$) for all conformally rigid graphs having less than 20 vertices. A case where this does not work is the Haar graph 565 on 20 vertices (see Fig. 11). However, there are other examples such as the non-Cayley vertex-transitive graph (24,23) on $n=24$ vertices, where all the other approaches failed while $cUU^{\top}$ indeed produces a valid certificate for $\lambda_{2}$.
Acknowledgment. SS was supported by the NSF (DMS-2123224). RT acknowledges helpful discussions with Sameer Agarwal.